An optical disk drive in your computer can spin a disk up to 10,000 rpm (about 1045 rad / s). If a particular disk is spun at 294.3 rad / s while it is being read, and then is allowed to come to rest over 0.234 seconds, what is the magnitude of the average angular acceleration of the disk? AND If the disk is 0.12 m in diameter, what is the magnitude of the linear acceleration of a point 1/8 of the way out from the center of the disk?
Answer #1

alfa = (w2-w1)/t

= (0 - 294.3)/0.234

= -1257.7 rad/s^2

------------------------

R = d/2 = 0.12/2 = 0.06 m

r = (1/8)*R = 0.06/8 = 0.0075

a_tan = r*alfa

= -0.0075*1257.7

= -9.43 m/s^2
Answer #2

I have solved this question earlier with different figures. Please workout using yours figures. If you need any further help just PM me. If I have helped you please rate me 5 stars first (before you rate anyone else)

An optical disk drive in your computer can spin a disk up to 10,000 rpm (about 1045 rad / s). If a particular disk is spun at 558.2 rad / s while it is being read, and then is allowed to come to rest over 0.435 seconds, what is the magnitude of the average angular acceleration of the disk?

____rad/s^2

If the disk is 0.12 m in diameter, what is the magnitude of the linear acceleration of a point 1/11 of the way out from the center of the disk?

______m/s^2

Answer

dw/t = alpha = 1283 rad/sec^2

alpha*r = a = 14 m/s^2